OpenStudy (anonymous):

40.)two and five -eights z plus 3/8 is equal is equal to 2-5z.. find z 41.)if number and -3 is divided by 10, the result is equal to the quotient obtained when the number minus two is divided by 3. what is the number??? 42.) the sum of a number and one-third of the number is equal to twice the number and 10. find the number??? 33.) what number is equal to its double? 39.) what number is equal to two-thirds of itself??

OpenStudy (anonymous):

This is not a call for help.. this is a call for answers. Please don't be lazy and try this on your own.

OpenStudy (anonymous):

excuse me but I did try them, and I don't understand them.. I wanted help with them, that was rude @love_jessika15

OpenStudy (anonymous):

40.)\[\left(2+\frac{5}{8}\right)z+\frac{3}{8}=2-5z \]Verify that\[z=\frac{13}{61} \]

OpenStudy (anonymous):

thanks I sorta get it

OpenStudy (anonymous):

not sure exzactly how u did it

OpenStudy (anonymous):

\[\left(2+\frac{5}{8}\right)z+\frac{3}{8}=2-5z \]\[8\left(\left(2+\frac{5}{8}\right)z+\frac{3}{8}\right)=(2-5z)8 \]\[3+21 z=16-40 z\]\[61 z=13\]Can you finish from here?

OpenStudy (anonymous):

I understood it now, I got z=4.69.. would u mind helping w/41 and 42 at least????? please if u don't care

OpenStudy (anonymous):

for 41.) I got 8 and 9/10 whatd u get @robtobey??????

OpenStudy (anonymous):

z=0.213115 or\[\frac{13}{61} \]

OpenStudy (anonymous):

are u on 41???

OpenStudy (anonymous):

Sorry. Have to break off for dinner now.

OpenStudy (anonymous):

wait, ur last message was that 40 r 41???????